Solve the matrix:

1 -1 2 1
-2 3 0 0
1 1 1 -4

(a, b, c, d) =

4
-4
3
Answer #1

solve the matrix I -1 2 -2 3 06-09) 1 1 1 are ; So, the 0 2 3 equations ark +20 +d4. -29 +36-4 a+b+c-4da 3. So there & variab2a- 4 t.c4d = 3-a 3 8 >>. 2a-4+36-12d39-30 120 = 13-59. > 30-12 we 2 8 get; 8-a 6/6 +3d & c 240 = 26 -10a t 27 d =-18 +99. q